# RONA: Pragmatically Diverse Image Captioning with Coherence Relations

### v1

@@ Abstract
Writing Assistants (e.g., Grammarly, Microsoft Copilot) traditionally generate diverse image captions by employing syntactic and semantic variations to describe image components. However, human-written captions prioritize conveying a central message alongside visual descriptions using pragmatic cues. To enhance pragmatic diversity, it is essential to explore alternative ways of communicating these messages in conjunction with visual content. To address this challenge, we propose RONA, a novel prompting strategy for Multi-modal Large Language Models (MLLM) that leverages Coherence Relations as an axis for variation. We demonstrate that RONA generates captions with better overall diversity and ground-truth alignment, compared to MLLM baselines across multiple domains. Our code is available at: https://github.com/aashish2000/RONA
@@ 1 Introduction
To address this challenge, in this work, we propose RONA (Relation-based cOhereNce-aware cAptioning), a novel prompting strategy for MLLMs inspired by the concept of Coherence Relations (CRs). Based on the principles of Discourse Theory, CRs provide a structured overview of image-text relationships Hobbs (1978); Kress (2009); O’Halloran et al. (2014). We evaluate their effectiveness in image captioning by using CRs as guidelines for generating captions that fulfill specific communicative functions while preserving semantic coherence.
@@
  • We propose RONA, a novel prompting strategy for MLLMs that leverages Coherence Relations (CRs) to generate pragmatically diverse image captions.

### v2

@@ Abstract
Writing Assistants (e.g., Grammarly, Microsoft Copilot) traditionally generate diverse image captions by employing syntactic and semantic variations to describe image components. However, human-written captions prioritize conveying a central message alongside visual descriptions using pragmatic cues. To enhance caption diversity, it is essential to explore alternative ways of communicating these messages in conjunction with visual content. We propose RONA, a novel prompting strategy for Multi-modal Large Language Models (MLLM) that leverages Coherence Relations as a controllable axis for pragmatic variations. We demonstrate that RONA generates captions with better overall diversity and ground-truth alignment, compared to MLLM baselines across multiple domains. Our code is available at: https://github.com/aashish2000/RONA
@@ 1 Introduction
To address this challenge, in this work, we propose RONA (Relation-based cOhereNce-aware cAptioning), a novel prompting strategy for MLLMs inspired by the concept of Coherence Relations (CRs). Based on the principles of Discourse Theory, CRs provide a structured overview of image-text relationships Hobbs (1978); Kress (2009); O’Halloran et al. (2014), modeling both contextual and pragmatics aspects of language Ma et al. (2025); Mavridou et al. (2015).
We evaluate their effectiveness in image captioning by using CRs as guidelines for generating captions that fulfill specific communicative functions while preserving semantic coherence.
@@
  • We propose RONA, a novel prompting strategy that leverages Coherence Relations (CRs) to generate pragmatically diverse image captions.
@@ 6 Conclusions
Acknowledgments
This research was in part supported by the U.S. National Science Foundation (NSF) award #1820609. Part of the research results were obtained using the computational resources provided by CloudBank (https://www.cloudbank.org/), which was supported by the NSF award #1925001.